OpenStudy (anonymous):

2x^3 - 8 How do I solve this?

hero (hero):

You can only solve an equation when it is equal to something. I see that you have posted an expression but what is it equal to?

OpenStudy (anonymous):

it's equal to zero

hero (hero):

2x^3 - 8 = 0 2x^3 = 8 x^3 = 8/2 x^3 = 4 x = cube root 4

OpenStudy (anonymous):

Oh I see! Thank You! : )